The price of a 2-hour plumbing repair job is $140 and the price of a 5-hour job is $290. Use the slope to predict the cost of a repair job lasting 4 hours.: The price of a 2-hour plumbing repair job is $140 and the price of a 5-hour job is $290. Use the slope to predict the cost of a repair job lasting 4 hours.
Answer by nerdybill(466)

The price of a 2-hour plumbing repair job is $140 and the price of a 5-hour job is $290. Use the slope to predict the cost of a repair job lasting 4 hours.
.
Think of it this way... The x-axis represents hours. The y-axis is cost of the job.
.
So, now we have two points:
(2, 140)
(5, 290)
.
Using "slope" formula of two points:
slope = (290-140)/(5-2) = 150/3
.
Now, using the slope (150/3) and one of the two points (2,140), we can derive the "slope - intercept" equation:
y = mx + b
plugging in what we know:
140 = (150/3)(2) + b
140 = (300/3) + b
140 = 100 + b
40 = b
.
Now we have the equation:
y = (150/3)x + 40
.
To predict the cost for 4 hour job, substitute 4 for x and solve for y:
y = (150/3)4 + 40
y = (600/3) + 40
y = 200 + 40
y = $240